OpenStudy (anonymous):

If a vector A⃗ has the following components, use trigonometry to find its magnitude and the counterclockwise angle it makes with the +x axis: Ax = 9.00lb , Ay = 5.00lb

OpenStudy (roadjester):

what is lb? pounds?

OpenStudy (samgrace):



OpenStudy (samgrace):

Magnitude is\[\sqrt{A_{x}^2 + A_{y}^2}\] Angle \[\tan \theta = A_{y}/A_{x} \] therefore \[\theta = \tan^{-1}(A_{y}/A_{x}) \]

OpenStudy (anonymous):

i plugged in the numbers for the last equation and it said the answer was wrong

OpenStudy (roadjester):

is your calculator in radians or degrees?

OpenStudy (anonymous):

in degreees

OpenStudy (roadjester):

29 degrees is incorrect?

OpenStudy (anonymous):

i found a different number but 29 was correct can you explain how you found that?

OpenStudy (roadjester):

\[\tan ^{-1}({\frac {A_y} {A_x}})= \tan ^{-1}({\frac {5} {9}})\]

OpenStudy (roadjester):

did your calculator give you 0.507?

OpenStudy (anonymous):

no this time i got 29 i must have entered it wrong the first time

OpenStudy (roadjester):

cool then

OpenStudy (anonymous):

thank you! if you have time one more thing, same type of problem but its with Ax = -24.0m/s , Ay = -34.0m/s i got 54.78 but it is supposedly the wrong answer

OpenStudy (roadjester):

ah, negatives are tricky

OpenStudy (anonymous):

yeah i just tried solving it the same and it wasn't working out i guess

OpenStudy (roadjester):

234.8

OpenStudy (anonymous):

thats correct, how did you get that

OpenStudy (roadjester):

your calculator gave you 54.8 right?

OpenStudy (anonymous):

right

OpenStudy (roadjester):



OpenStudy (roadjester):

54.8 is also correct

OpenStudy (roadjester):

divide Ay/Ax and tell me what you get then take tangent of BOTH angles you got

OpenStudy (roadjester):

you'll notice there's no difference

OpenStudy (anonymous):

well i did -34/-24 and then i took the \[\tan^{-1} \] 1.41667 and got 54.78 what would the other angle be?

OpenStudy (roadjester):

oops, sorry, i meant tan(54.78) tan(234.78) don't do \[tan^{-1}\]

OpenStudy (anonymous):

ok yeah i get the same angles. but how did you get the 234.78?

OpenStudy (roadjester):

you have to do that yourself. see, while a calculator is good at performing hard calculations, it's not very smart. so what it does is, it goes around the 4 quadrants starting from quadrant 1 until it finds an angle that will give you 1.41667. So it never reaches 278 because 54.78 already gives you that number, does that make sense?

OpenStudy (roadjester):

so for example, sin(30)=sin(150)=1/2 but if you do sin ^-1 (1/2) you will NEVER get 150, you will always get 30. does that make sense?

OpenStudy (anonymous):

yes that makes sense. so how do i do it myself?

OpenStudy (roadjester):

ok, that really depends on the trig function you're talking about. for tangent, just add 180 but for sin, it has to be 180- the angle.

OpenStudy (roadjester):

so 180-30=150

OpenStudy (roadjester):

54.78+180=278.78

OpenStudy (roadjester):

but just to check, you can always just do the reverse after adding or subtracting. there is a derivation to all of this

OpenStudy (anonymous):

ok that makes total sense thank you! this has been very helpful

OpenStudy (roadjester):

hey no problem

OpenStudy (roadjester):

and remember

OpenStudy (roadjester):

calculators are stupid

OpenStudy (roadjester):

=D

OpenStudy (roadjester):

sorry, just had to say that

OpenStudy (anonymous):

haha no problem they cause me more harm then good most of the time